OpenStudy (anonymous):

Solve for x: 1 over 2 times x equals 3 and 7 over 15

OpenStudy (oaktree):

Do you mean\[\frac{ 1 }{ 2 }x = 3 \frac{ 7 }{ 15 }\]

OpenStudy (oaktree):

Or is the x in the denominator?

OpenStudy (anonymous):

No its not the denominator

OpenStudy (oaktree):

Okay. So we want to start by isolating the x. But mixed fractions are gross, so first we want to make it an improper fraction to make our lives easier. I'm assuming you know how to turn a mixed fraction into an improper fraction?

OpenStudy (anonymous):

Yes

OpenStudy (oaktree):

Okay. So what would the mixed fraction be as an improper fraction?

OpenStudy (oaktree):

And thanks for the medal, by the way.

OpenStudy (anonymous):

52 over 1552 over 15

OpenStudy (oaktree):

Wow. That came out jumbled! But yeah, 52/15.

OpenStudy (anonymous):

Omg ehasorry this is a tablet

OpenStudy (oaktree):

No worries. So now to isolate the x. In essence, multiplying by one-half is the same as dividing by 2. So if we want to isolate the x, what should we multiply the left side by?

OpenStudy (anonymous):

The recripocal

OpenStudy (oaktree):

Which is?

OpenStudy (anonymous):

Oops typo...

OpenStudy (oaktree):

That's fine. But what's the reciprocal of 1/2?

OpenStudy (anonymous):

Two over one

OpenStudy (oaktree):

Also known as...?

OpenStudy (anonymous):

Brb I'm movie to a desktop

OpenStudy (oaktree):

That's okay. Anyways, that's just 2. So we multiply both sides by 2. We get just x on the left hand side, and we get \[\frac{ 52 }{ 15 } \times 2\] Which is the same as \[\frac{ 104 }{ 15 }\]Which, if you need it as a mixed fraction, is the same as \[6 \frac{ 14 }{ 15 }\]

OpenStudy (anonymous):

Thank you

OpenStudy (oaktree):

No problem!